The activation energy of a particular reaction is 83.1 kJ/mol. By what factor will the rate constant increase when the temperature is increased from 50 °C to 60 °C?
Answer #1


Ti =50 °C =(50 + 273)K =323K T2 = 60 °C = (60+ 273)K = 333K Activation energy, Ea=83.1x10° J/mol Use the following equation t